QUESTION 2 (15 Marks) Dagbreek CC sells coffee beans. The company have several outlets in shopping malls across the Western Cape. However over the last 5 years average costs increased by more than 10% per annum driven primary by rental increases on rental space in the malls. As part of a revised business strategy, the CEO is considering selling beans to corporate offices around Cape Town, hoping to increase revenue. The company is considering hiring 12 sales representatives who are constantly on the road driving coffee bean sales to offices around the city The CEO of the company is considering two plans for paying his 12 sales people. The variable cost plan is based on a straight commission of R 48 per kilogram sold, and the mixed cost plan is based upon a salary of R24,500 plus a commission of R 23,50 per kilogram sold. Unit variable costs are R 210 (excluding any commission) and the fixed expenses are estimated at R 850,000 (excluding all sales people salaries)

REQUIRED Compute the point of indifference (i.e where the operating income under both plans will be indentical).

At a projected sales volume of 2,500 kilograms, which plans should Dagbreek adopt and why?

You have a meeting with the CEO. Explain to him the difference between variable costing & absorption costing.
Answer #1
Units sold 2500
Fixed Cost 850000
Variable Cost 210 p.unit i.e. 525000 (210*2500)
So, Total Cost 850000+525000 =1375000
Now there are 2 plans
1) or 2)
Variable Commission = 48 per KG Variable Commission = 23.5 per KG plus
Fix salary 24500 *12 = 294000
i.e. 48 * 2500 = 120000 i.e. 23.50 * 2500 + 294000 = 352750
in Plan 1 commission payable is less than plan 2, therefor for 2500 units Dagbreek should adopt it
Now, Calculation of point of indifference
Total Fix salary is R 294000 and difference of variable commission is 48 - 23.50 = 24.50
To calculate point of indifference formula is = Fixed Salary / Difference of variable commission
= 294000 / 24.50
= 12000
So, up to 12000 units Plan 1 of variable commission will cost less and after 12000 units of sale Plan 2 will be cost efficient

Variable costing is defined as an accounting method for production expenses where only variable costs are included in the product cost.

Absorption costing includes all costs associated with a production process that is assigned to the units produced